OpenStudy (anonymous):

How do you find the value of b? The line through the points (-5, -5) and (4,b) with the slope - 1/4.

OpenStudy (anonymous):

well since you know that slope is \[\frac{ \Delta y }{ \Delta x }\] you know that for every y it goes down it goes over four x

OpenStudy (anonymous):

so you reverse this since you are going from a negative x to a positive x so for every 4 you go over you go up one x

OpenStudy (anonymous):

so -5+4=-1 and then up one on the y -5+1=-4 so you next point is (-1,-4)

OpenStudy (anonymous):

so you do this until you reach the value of 4 for you x value

OpenStudy (anonymous):

so what do you think b is?

OpenStudy (anonymous):

I wasn't sure because my equation was -1(b+5)/9 = 9(-1/4)

OpenStudy (anonymous):

yeah I have no clue where that equation came from it is just simpler to use the method I just described

OpenStudy (anonymous):

I don't quite understand your method

OpenStudy (anonymous):

so you know that slope is change in y over change in x right

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

since i tend to misplace values, I just move the points ... place one of them at the origin (-5, -5) and (4 , b) +5 +5 +5 +5 ----------------- 0 , 0 9, b+5 ^^^^^ i can use this to define the slope, y/x slope = (b+5)/9 and equate that with the given slope to solve for b

OpenStudy (anonymous):

I got the answer -29/4

OpenStudy (amistre64):

yes